OpenStudy (anonymous):

Solve and check: 1= 7/a-2 - 28/a^2-4 (fractions) a= ??

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

\[1=\frac{7}{x-2}-\frac{28}{x^2-4}\] factor and add gives \[1=\frac{7}{x-2}-\frac{28}{(x-2)(x+2)}=\frac{7(x+2)}{(x+2)(x-2)}+\frac{28}{(x-2)(x+2)}\]

OpenStudy (saifoo.khan):

Satellite to the rescue again!

OpenStudy (anonymous):

\[\frac{7x+42}{(x-2)(x+2)}=1\] \[7x+42=x^2-4\] \[x^2-7x+46=0\]

OpenStudy (anonymous):

a = - 2 l + 7 / l

OpenStudy (anonymous):

Or was that a 1. Sorry,

OpenStudy (anonymous):

no real solution damn

OpenStudy (anonymous):

oh no i made a typo \[x^2-7x-46=0\]

OpenStudy (saifoo.khan):

Damn^

OpenStudy (anonymous):

a =5?

OpenStudy (anonymous):

i need quadratic formula for this one. did i make an error?

OpenStudy (anonymous):

i messed up some where. let me see where

OpenStudy (anonymous):

bush league error i made

OpenStudy (anonymous):

I think it's 5.

OpenStudy (anonymous):

\[1=\frac{7}{x-2}-\frac{28}{(x-2)(x+2)}=\frac{7(x+2)}{(x+2)(x-2)}-\frac{28}{(x-2)(x+2)}\]

OpenStudy (anonymous):

i dropped the minus sign!

OpenStudy (saifoo.khan):

Damn^

OpenStudy (anonymous):

\[7x-14=x^2-4\] \[x^2-7x+10=0\] \[(x-2)(x-5)-0\] so \[x=5\] because \[x=2\] is not allowed (divide by zero and all)

OpenStudy (anonymous):

Did you get 5 too?

OpenStudy (anonymous):

oh lol just saw this. cool.